Journalist Report – March 6th

Journalist report, 03/06/18: CSI: MDRS

An unfathomable mystery arose in the Hab in the early morning. Benoît and Jérémy had cooked some popcorn chicken yesterday, yet Victoria did not want to eat her part on the evening, and spared it for today in her room. However, as she woke up, all the chicken was gone! Several people were suspected: mice at first, yet there were no crumbs around the bowl. Jérémy and Benoît then, as the first one was not in the Hab when she revealed her loss, and as the lack of food for the second is no longer a secret! However, Victoria would have certainly heard it if someone broke into her room to steal her chicken. The case is still open!

Today was quite a special day in the Hab, as we received two journalists from one of the main French television stations TF1. Presenter Axel Monnier and cameraman Bertrand Guez arrived short after 7am, right in the middle of our sport session. They could thus observe us sweat during our last burpees, such a nice start. After a small breakfast with a piece of French bread (well, rather bread with the French flag in it), it was time for most of us to go on EVA! Our both journalists were part of the mission: Axel even followed us in putting a spacesuit on! Gabriel, Jérémy, Victoria and I as their leader were completing the quite large team.

As it has now become routine, we made our first stop at Pooh’s corner to put the LOAC back in its place. It will yet not stay there long, as Jérémy wants to have the time to complete some measurements inside the Hab as well. We then went for the road of Lith Canyon. It seemed that Bertand sometimes needed a specific framework, as he liked to put himself in acrobatic positions on the rover, in order to be able to record us following the rover on our ATV. However, the road was rather smooth, and we arrived at Lith Canyon without any problem.

Our guests were glad to visit this place, which is probably the nearest to a Martian landscape we could get here. The walk was really pleasant, yet sometimes maybe a bit harsh for Bertrand, who had to bear his 30-pound 50000-dollard camera on his shoulder. As we reached another cliff, we decided to stop our way down: Axel and Bertrand could thus deploy their drones and take some aerial shots, after a short interview of the four of us. It was yet already time to go back to the Hab, where our journalists wanted to take some final shots.

They could question us deeper during our lunch, as they wanted to know what our thoughts about that kind of mission were. We finally could show them some of the experiments we had led for the last weeks, as LOAC and MegaARES’ ones, Laurent’s solar panels or Jérémy’s human factors one. But as the afternoon reached its end, it was already time for them to leave us, ready to reach Salt Lake City where another report awaited them. We still have a lot to do in the Hab before the end of the mission, especially with our rotation video, as we still have many shots to do.

Operations Report – March 6th

Crew 189 Operations Report 6March2018

SOL: 16

Name of person filing report: Gabriel PAYEN

Non-nominal systems: –

Notes on non-nominal systems: –

Generator (hours run): Turned off at 09:AM and turned on at 6:20PM

Solar— SOC 72 % (Before generator is run at night)

Diesel – 47 %

Propane – 21 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 9.5 gallons.

Water (trailer) – 0 gallons.

Water (static) – 500 (90%) gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 131 724.2

Toilet tank emptied: NO

ATV’s Used: 350.2 & 350.3

Oil Added : NO

ATV Fuel Used: Gals

Hours the ATVs were Used today: 3

Notes on ATVs: –

Deimos rover used: YES

Hours: Start : 112.7 , End : 113.8

Beginning charge: 100%

Ending charge: 86%

Currently charging: YES

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: YES

Hours: Start : 21.9 , End : 23

Beginning charge: 100%

Ending charge: 70%

Currently charging: YES

Opportunity rover used: NO


Beginning charge: 100

Ending charge: –

Currently charging: NO

Curiosity rover used: NO

Hours: 18

Beginning charge: 100%

Ending charge: –

Currently charging: NO

HabCar used and why, where : used to get aquaponics supplies and packages from crew 190

General notes and comments: –

Summary of internet: –

Summary of suits and radios: –

Summary of Hab operations: –

Summary of GreenHab operations:

Summary of ScienceDome operations: –

Summary of RAMM operations:-

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Greenhab Report – March 6th

GreenHab Report
6 March 2018

Environmental control:

(door closed)
Heating once the sun was down
80 % Shade cloth on

Average temperatures:
Around 25 °C at 9:00 am, around 26 °C the rest of the day

Hours of supplemental light: none

Changes to crops: none

Daily water usage for crops: around 9 gallons

Time(s) of watering for crops: 9:00 am and 7pm

Narrative: none

Support/supplies needed: none

Commander Report – March 6th

Sol 16

Title: Two new crewmembers!

Dear Earth,

Today, two French journalists from French channel television (TF1) joined the crew on the Red Planet for our typical tasks: physical training, breakfast, morning EVA to check our experiments and explore the Martian surface, engineering check, cooking, EVA debriefing, experiments in the Hab and the Science Dome…

They left our Martian habitat at the end of the afternoon, really happy of what they discovered during their trip! Once again, we were happy to see new faces after almost three weeks confined with the same people. We are so excited to see the results of this TV report and to watch it on the main French TV channel! We were really glad because they asked us many questions, and very interesting questions!

After our guests left, we were once again all busy in the Hab: we did the human factors experiment of Gabriel (the KTNE game), Alex worked on the video presenting our rotation, and I worked on my last experiments.

Moreover, today was a big day, not only because we had guests, but also because we changed the teams of the KTNE experiment: it is funny to see that we have difficulties when we have to discover new partners while playing and defusing bombs.

We have three Sols left on this very nice planet, but we still have a lot of work to do on our experiments before leaving. Luckily, we will have our EVA tomorrow afternoon so perhaps we will sleep more in the morning after these two early wake up!

Ad Astra!

Sol Summary – March 6th

Crew 189 Sol Summary Report 06 Mar 2018

Sol 16

Summary Title: TV stars

Author: Louis Mangin

Mission Status: All nominal

Sol Activity Summary:
Today was another special day: a journalist duo came from Washington
to shoot a report for the news on a French national TV. They were very
curious and friendly and should have taken great shots of Lith Canyon
and the station. The reporter followed us wearing a spacesuit during
the EVA. Let’s hope the result looks great!

The afternoon was quieter: human factors experiments, astronomy,
experiments, video editing, demonstrations/interviews for the

Look Ahead Plan: Emergency procedures training vol. 2 in the morning,
small EVA in the afternoon to deploy Solar Panels, GPS guidance
system, shoot missing footage for the rotation video

Anomalies in work: –

Weather: Sunny.

Crew Physical Status: Fine.

EVA: Lith Canyon + Battery change

Reports to be filed:
Greenhab Officer
EVA report
EVA request
Astronomy report

Support Requested: –

EVA Report – March 6th

EVA Report:

EVA #12 Crew members: Alexandre Martin (EVA leader), Gabriel Payen, Jérémy AUCLAIR, Victoria Da Poian, Axel Monnier (journalist), Bertrand Guez (journalist)

Author: Alexandre Martin


Near Pooh’s Corner: 12S-518900E-4250900N

Lith Canyon: 12S-518300E-4256500N

Time: departure at 9:30 a.m. Duration: 3h


  • Replace the LOAC
  • Explore Lith Canyon


We made some shots with the journalist in front of the Hab, and then we left the Hab to replace the LOAC at Pooh’s corner. We went to Lith Canyon to show our journalists a Martian-type environment. They took some shots with their drone, and we came back at the Hab, where they took some final shots.

9:30 am: EVA#12 left airlock

9:45 am: EVA#12 left the Hab

9:46 am: EVA#12 arrived at Pooh’s corner

10:05 am: EVA#12 left Pooh’s Corner

10:30 am: EVA#12 arrived at Lith Canyon

11:50 am: EVA#12 left Lith Canyon

12:15 pm: EVA#12 back at the Hab

12:30 pm: EVA#12 back in the Hab

EVA Report – March 5th


Here is today’s EVA report also comming from Jérémy:

EVA Report:
EVA #11 Crew members: Jérémy Auclair (EVA leader), Louis Mangin, Gabriel Payen, Benoit Floquet

– Around the Hab
– Near Pooh’s Corner: 12S-518900E-4252000N
– Near Mid Ridge Planitia: 12S-517500E-4251500N

Time: departure at 09:12 a.m. Duration: 2 hour 24 minutes

When describing shortly today’s EVA, we could say that concerning our experiments it was a failure. After arriving at the instruments’ position near Pooh’s corner, we found out that both batteries were empty. Therefore, we plugged the new battery on MegaARES to start transferring data on Gabriel’s USB key. We took one empty battery (to start charging it) and went back to the Hab to start Benoit’s localization experiment. After more than five minutes the GPS chip still hadn’t found any signal.

We decided to continue the walk to get on the Hab ridge. We took quite a few pictures, Louis with his camera and me with mine, on which I had mounted my orange filter (for my personal pictures). After we arrived over the Hab we headed back to the Hab. We took the ATVs to go back to Pooh’s Corner to finish MegaARES’ data transfer.

Once the transfer was done we plugged the LOAC on the battery. But once again the power system was broken. This time it was the switch which was broken. We took the decision to bring the two batteries and the LOAC back to the Hab to fix the power system.

Unfortunately, on the way back to the Hab, one of the batteries fell of one of the ATV 30 meters away of the Hab because of a bump on the road. It broke and started to leak slightly. We must have attached it too weakly because we had to bring back two batteries instead of one. I quickly asked Habcom to put a garbage bag in the engineering Airlock. 3 Minutes later we put the leaking battery in the plastic bag to confine it. Once this (small) crisis was taken care of we started the final engineering check and ended the EVA.

– LOAC: Test of the battery level
– MegaARES: Replacement of the battery
– Localization experiment: test with increasing uncertainty on the positioning measure

Vehicles: 4 ATVs

9:12 a.m.: EVA#11 left Airlock
9:20 a.m.: EVA left the Hab using the four ATVs
9:33 a.m.: Arrival on the working place and beginning of battery tests and data transfert
9:39 a.m.: Return to the ATVs
9:42 a.m.: Turning on the ATVs to go back to the Hab
9:46 a.m.: Arrival to the Hab, Benoit starts his GPS instrument 9:50 a.m.: Beginning of Walk to find a way to climb on the ridge 9:57 a.m.: Benoit’s instrument still doesn’t work
10:22 a.m.: Louis finds a safe way to climb on the ridge
10:32 a.m.: Arrival on top of the ridge
10:42 a.m.: Beginning of way down the ridge
11:00 a.m.: Departure to the instruments on Pooh’s Corner
11:18 a.m.: Departure to the Hab with both batteries and the LOAC 11:36 a.m.: Entry in the Main Airlock



Greenhab Report – March 5th

Hello again Capcom,

Here is today’s GreenHab report coming from Jérémy:

GreenHab Report
5th March 2018

Environmental control:

(door closed)
Heating once the sun was down
80 % Shade cloth on

Average temperatures:
Around 30 °C at 11:30 am, around 23 °C the rest of the day

Hours of supplemental light: none

Changes to crops: 5 cherry tomatoes harvested

Daily water usage for crops: around 9 gallons

Time(s) of watering for crops: 10:00 am and 9pm

Narrative: none

Support/supplies needed: none



Copyright © The Mars Society. All rights reserved. | Main Site